Circular No. 8262

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003lq IN UGC 5 Further to IAUC 8229, H. Pugh and R. J. Foley report the LOSS

discovery, on KAIT images taken on 2003 Dec. 28.1 (mag ∼ 17.2) and 2004 Jan. 6.1 UT (mag ∼ 15.9), of an apparent supernova located at α = 0^{h}03^{m}04^{s}.02, δ = −1^{o}54′45′′.5 (equinox 2000.0), which is 24′′.5 west and 5′′.6 north of the nucleus of UGC 5. A KAIT image taken on 2003 Nov. 21.2 showed nothing at this position (limiting mag ∼ 19.0).

NOVAE IN M31 Some additional apparent novae in M31 have been reported:

2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
Dec. 18.767 0^{h}43^{m}04.^{s}99 +41^{o}12′21′′.7 16.7 240′′ E, 215′′ S
Dec. 20.708 0 42 53.22 +41 22 35.5 16.1 100′′ E, 387′′ N
Dec. 20.708 0 42 54.08 +41 15 12.1 16.7 110′′ E, 56′′ S
Dec. 20.708 0 43 04.77 +41 12 20.8 18.5: 231′′ E, 228′′ S

The first object above was reported by M. Mobberley (Cockfield, England, from unfiltered CCD images; via G. M. Hurst), whose data are tabulated; by O. Smirnova (from blue photographs taken by A. Barzdis on 2003 Dec. 19.696 and 19.726 UT with the Schmidt telescope of the Astrophysical Observatory at Baldone; via A. Alksnis, University of Latvia; position end figures 04^{s}.8, 23′′); and by K. Hornoch (cf. IAUC 8248; R-band CCD frames taken on Dec. 23.695 and 23.985; position end figures 04^{s}.77, 23′′.0; offset 231′′ east and 226′′ south of the galaxy center). The other three objects were found by M. Fiaschi et al. (cf. IAUC 8238) on Hα images. Further magnitudes for the first object (unfiltered CCD, with A = M. Armstrong, Rolvenden, England, except H = Hornoch, R-band; and S = Smirnova et al., blue photo): 2002 Sept. 28, [20.0 (T. Boles, Coddenham, England); 2003 Aug. 8, [20.0 (A); Dec. 9.732, [19.7 (H); 12, [19.5 (S); 17.895, 18 (A); 19.696, 18.6 (S); 23.695, 17.8 (H); 23.985, 17.9 (H); 24.825, 19.0 (H). Further magnitudes for the second object tabulated above (F = Fiaschi et al., Hα): Dec. 13, [18.5 (F); 23.695, R = 18.0 (H); 23.726, 15.7 (F); 24.825, R = 18.1 (H). Further magnitudes for the third object: Dec. 13, [18.5 (F); 23.695, 18.3 (H); 23.726, 16.4 (F); 24.825, 18.2 (H). Further Hα magnitudes for the fourth object: Dec. 13, [18.5 (F); 23.726, 17.9 (F). Hornoch’s magnitudes for the Fiaschi-et-al. object on IAUC 8253 and 8254: Nov. 29.869, 17.9 (near edge of field); Dec. 7.678, 17.6; 8.833, 17.9; 9.732, 17.7; 24.825, 18.9.

2004 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green